# AnyAttack: Towards Large-scale Self-supervised Generation of Targeted Adversarial Examples for Vision-Language Models

### v1

@@ 4 Experiments
Attack Method MSCOCO TR@1 TR@5 TR@10 IR@1 IR@5 IR@10 R@Mean ViT-B/16 AttackVLM-ii 0.4 1.0 1.4 0.24 1.08 2.16 1.05 AttackVLM-it 0.2 1.4 1.8 0.16 1.16 2.12 1.14 SASD-WS-Cos 6.0 17.0 24.8 9.08 24.39 34.55 19.30 SASD-WS-MSE 4.8 18.4 25.6 8.20 25.87 35.15 19.67 SU-Cos 6.8 20.4 27.8 11.11 25.70 33.34 20.86 SU-MSE 6.8 20.6 27.0 10.83 25.10 32.62 20.49 AnyAttack-Cos 8.6 21.2 29.6 10.80 27.59 37.50 22.55 AnyAttack-Bi 12.2 26.2 33.8 12.63 31.71 40.86 26.23 AnyAttack-Cos w/ Aux 8.4 24.8 33.0 11.59 32.10 44.98 25.81 AnyAttack-Bi w/ Aux 14.8 36.8 48.0 17.59 42.02 56.05 35.88 ViT-L/14 AttackVLM-ii 0.2 1.0 1.6 0.24 0.60 1.32 0.83 AttackVLM-it 0.4 0.8 1.4 0.12 0.76 1.48 0.83 SASD-WS-Cos 3.8 11.6 18.8 7.20 18.43 26.35 14.36 SASD-WS-MSE 5.4 14.6 20.6 6.00 18.23 26.47 15.22 SU-Cos 3.0 10.4 13.2 6.19 14.99 20.07 11.31 SU-MSE 3.4 11.2 17.4 6.63 15.27 19.79 12.28 AnyAttack-Cos 3.8 14.0 22.8 7.36 20.71 27.55 16.04 AnyAttack-Bi 4.8 16.0 23.6 8.20 22.31 29.11 17.34 AnyAttack-Cos w/ Aux 9.4 24.6 37.0 11.51 32.62 48.18 27.22 AnyAttack-Bi w/ Aux 12.0 34.0 47.4 15.67 39.34 53.54 33.66 ViT-L/14 $\times$ 336 AttackVLM-ii 0.2 0.6 1.6 0.16 1.12 2.04 0.95 AttackVLM-it 0.2 0.6 1.8 0.32 0.96 1.76 0.94 SASD-WS-Cos 2.8 10.8 16.4 6.52 18.31 26.19 13.50 SASD-WS-MSE 4.4 13.6 19.2 6.72 18.23 25.71 14.64 SU-Cos 2.4 8.0 11.2 4.88 13.79 18.39 9.78 SU-MSE 3.6 8.2 13.2 6.40 14.51 19.19 10.85 AnyAttack-Cos 4.6 11.0 16.6 5.96 17.67 24.23 13.34 AnyAttack-Bi 3.6 14.4 19.0 7.64 19.79 26.83 15.21 AnyAttack-Cos w/ Aux 9.0 23.2 37.2 11.68 34.03 47.62 27.12 AnyAttack-Bi w/ Aux 12.0 33.2 46.8 14.79 39.22 53.06 33.18
@@ 4.1 Experimental Setup
We first employed the state-of-the-art (SOTA) targeted adversarial attack for VLMs, referred to as AttackVLM (Zhao et al., 2024). This method includes two variations: AttackVLM-ii and AttackVLM-it, which are based on different attack objectives. Both methods utilize the CLIP ViT-B/32 image encoder as the surrogate model, consistent with our approach. Additionally, we incorporated two targeted adversarial attacks designed for visual classification models: SU (Wei et al., 2023) and SASD-WS (Wu et al., 2024). Since the original cross-entropy loss used in these methods is not suitable for vision-language tasks, we modified them to employ cosine loss and mean squared error (MSE) loss to match targeted images. These modified methods are denoted as SU-Cos/SASD-WS-Cos and SU-MSE/SASD-WS-MSE, respectively. For the SU attack, the surrogate model is configurable, and we set it to align with our proposed method, namely, CLIP ViT-B/32. For the SASD-WS attack, we utilized the officially released weights, as its surrogate model includes a self-enhancement component. We denote our proposed methods as AnyAttack-Cos, AnyAttack-Bi, AnyAttack-Cos w/ Aux, and AnyAttack-Bi w/ Aux. These represent AnyAttack fine-tuned with $\mathcal{L}_{\text{Cos}}$, fine-tuned with $\mathcal{L}_{\text{Bi}}$, fine-tuned with $\mathcal{L}_{\text{Cos}}$ using auxiliary models, and fine-tuned with $\mathcal{L}_{\text{Bi}}$ using auxiliary models, respectively.
Datasets, Models, and Tasks
@@
For the downstream datasets, we utilize the MSCOCO, Flickr30K, and SNLI-VE datasets. We employ a variety of target models, including CLIP, BLIP, BLIP2, InstructBLIP, and MiniGPT-4. The downstream tasks we focus on are image-text retrieval, multimodal classification, and image captioning. For each task, we selected the top 1,000 images. Additionally, following the methodology outlined in (Zhao et al., 2024), we used the top 1,000 images from the ImageNet-1K validation set as clean (unrelated) images to generate adversarial examples.
Metric
@@
In this work, we examine perturbations constrained by the $\ell_{\infty}$ norm, ensuring that the perturbation $\delta$ satisfies the condition $\|\delta\|_{\infty}\leq\epsilon$, where $\epsilon$ defines the maximum allowable magnitude of the perturbation. We use the attack success rate (ASR) as the primary evaluation metric to assess the performance of targeted adversarial attacks. The calculation of ASR varies slightly depending on the specific task. For instance, in image-text retrieval tasks, ASR represents the recall rate between adversarial images and their corresponding ground-truth text descriptions. In multimodal classification tasks, ASR refers to the accuracy of correctly classifying pairs of ”adversarial image and ground-truth description.” Essentially, ASR is calculated by replacing clean images with their adversarial counterparts and then computing the relevant task-specific evaluation metric.
Implementation Details
@@
We pre-trained the decoder for 520,000 steps on the LAION-400M dataset (Schuhmann et al., 2021), using a batch size of 600 per GPU on three NVIDIA A100 80GB GPUs. The optimizer used was AdamW, with an initial learning rate of $1\times 10^{-4}$, which was adjusted using cosine annealing. For the downstream datasets, we fine-tuned the decoder for 20 epochs using the same optimizer, initial learning rate, and cosine annealing schedule. We deployed two auxiliary models: the first is a ViT-B/16 trained from scratch on ImageNet-1K, utilizing official weights from torchvision, and the second is the EVA model (Fang et al., 2023, 2024), a ViT-L/14 model that employs masked image modeling and has been fine-tuned on ImageNet-1K. The maximum perturbation $\epsilon$ for $\delta$ was set to $\frac{16}{255}$, and $K$ was set to 5. In the pre-training stage, the initial temperature $\tau_{0}$ was set to 1, the final temperature $\tau_{\text{final}}$ was set to 0.07, and the total steps $T$ were set to 10,000. In the fine-tuning stage, the temperature $\tau$ was fixed at 0.07. More details can be found in the Appendix.
@@ 4.2 Evaluation on Image-Text Retrieval
In this subsection, we compare the performance of our method against baseline approaches on the image-text retrieval task. Table 2 presents the results on the MSCOCO dataset, while results on the Flickr30K dataset are detailed in the Appendix. The following key observations can be made:
@@
  • Advantages of Bidirectional InfoNCE Loss: The bidirectional InfoNCE loss $\mathcal{L}_{\text{Bi}}$ shows clear advantages for retrieval tasks, with AnyAttack-Bi consistently outperforming AnyAttack-Cos.
@@ 4.3 Evaluation on Multimodal Classification
Here, we compare the performance of our attack with the baselines on the multimodal classification task. Table 3 presents the results on the SNLI-VE dataset. Our method, AnyAttack-Cos w/ Auxiliary, achieves the highest performance, surpassing the strongest baseline, SASD-WS-MSE, by 20.0%. This underscores the effectiveness of our attack in multimodal classification tasks.
@@ 4.4 Evaluation on Image Captioning
Here, we evaluate the performance of our attack on the image captioning task using the MSCOCO dataset. The VLMs take adversarial images as input and generate text descriptions, which are then assessed against the ground-truth captions using standard metrics. Table 4 presents the results across four VLMs: InstructBLIP, BLIP2, BLIP, and MiniGPT-4. Our attack AnyAttack-Cos w/ Auxiliary consistently demonstrates superior performance across all evaluation metrics, outperforming the baseline attacks on each VLM.
@@ 4.5 Transfer to Commercial VLMs
Figure 3: Example responses from 3 commercial VLMs to targeted attacks generated by our method.
Here, we transfer the targeted adversarial images generated by the pre-trained decoder (referred to as AnyAttack-Pre) to three commercial VLMs, including Claude’s Sonnet, Microsoft’s Copilot, and Google’s Gemini. We utilized the publicly available web interfaces of these models. Figure 3 illustrates the example responses of the three commercial models, with more examples are provided in the Appendix. No prior conversation context or constraints were applied; the only prompt used was ”Describe this image.” The portions of the VLM responses highlighted in red correspond to the target images, showcasing the effectiveness of our attack.
Figure 4: Performance comparison between different configurations of AnyAttack for the image-text retrieval task on MSCOCO. The plot shows the comparative performance of models initialized from scratch, pre-trained, and fine-tuned, alongside the impact of auxiliary models and different fine-tuning objectives on retrieval tasks.
@@ 4.6 Further Analysis
Ablation Study
We perform an ablation study on the MSCOCO dataset for the image-text retrieval task to evaluate the impact of three key components in our approach: 1) Training approach: Pre-trained, fine-tuned, or trained from scratch. 2) Auxiliary models: With or without auxiliary model integration. 3) Fine-tuning objective: Cosine similarity loss ($\mathcal{L}_{\text{Cos}}$) vs. bidirectional contrastive loss ($\mathcal{L}_{\text{Bi}}$).
The results, summarized in Figure 4, reveal the following: 1) Training approach: Fine-tuning a pre-trained model achieves the highest performance, while training from scratch yields significantly worse results, indicating that pre-training is critical for task adaptation.
Figure 5: Efficiency Analysis: Memory Usage vs Time Consumption.
2) Auxiliary models: The inclusion of auxiliary models consistently improves performance, highlighting their role in enhancing transferability. 3) Fine-tuning objective: The bidirectional contrastive loss ($\mathcal{L}_{\text{Bi}}$) consistently outperforms the cosine similarity loss ($\mathcal{L}_{\text{Cos}}$), demonstrating its effectiveness in improving the alignment of image and text embeddings.
Efficiency Analysis
@@ 5 Conclusion
In this paper, we introduced AnyAttack, a novel self-supervised framework for generating targeted adversarial attacks on VLMs. Our approach overcomes the scalability limitations of previous methods by enabling the use of any image to serve as a target for attack target without label supervision. Through extensive experiments, we demonstrated the effectiveness of AnyAttack across multiple VLMs and vision-language tasks, revealing significant vulnerabilities in state-of-the-art models. Notably, our method showed considerable transferability, even to commercial VLMs, highlighting the broad implications of our findings.
@@
Appendix A Appendix
In this appendix, we describe more implementation details and additional experiment results.
A.1 Implementation Details
Table 5: Configuration details of VLMs.
Model Configuration
BLIP2 caption_coco_opt2.7b
InstructBLIP vicuna-7b-v1.1
Mini-GPT4 minigpt4_llama2_7b
In this section, we provide additional details regarding the experimental setup. For AttackVLM, we utilized the official code [^2] https://github.com/yunqing-me/AttackVLM., while for SU and SASD-WS, we employed the TransferAttack tool [^3] https://github.com/Trustworthy-AI-Group/TransferAttack.. For VLMs integrated with LLMs, we used the LAVIS library [^4] https://github.com/salesforce/LAVIS. and the MiniGPT-4 repository [^5] https://github.com/Vision-CAIR/MiniGPT-4.. More details are provided in Table 5. Regarding MiniGPT-4, it tends to generate detailed responses, even when the prompt “Describe this image in one short sentence only”. Occasionally, it outputs multiple sentences, which affects its scoring in the image captioning task.
A.2 Additional Experiment Results
Additional Results on Image-text Retrieval
Table 6: Comparison of retrieval performance on the Flickr30k dataset. TR@1, TR@5, and TR@10 represent text retrieval, while IR@1, IR@5, and IR@10 represent image retrieval. R@Mean is the average of all retrieval metrics. Our proposed methods are italicized, the best results are highlighted in bold, and the second-best results are underlined.
Method Flickr30k TR@1 TR@5 TR@10 IR@1 IR@5 IR@10 R@Mean ViT-B/16 AttackVLM-ii 0.2 1.8 2.4 1.00 2.00 2.80 1.70 AttackVLM-it 0.4 1.6 2.6 0.20 1.60 3.00 1.57 SASD-WS-Cos 4.4 12.6 19.8 7.00 17.20 25.20 14.37 SASD-WS-MSE 4.2 12.2 18.0 5.20 16.00 23.00 13.10 SU-Cos 5.0 17.2 25.2 8.40 20.20 27.00 17.17 SU-MSE 5.8 15.8 24.2 9.40 21.60 30.20 17.83 AnyAttack-Cos 7.0 18.8 24.8 11.40 25.40 33.00 20.07 AnyAttack-Bi 8.8 22.4 30.0 11.80 27.20 37.20 22.90 AnyAttack-Cos w/ Aux 8.8 24.4 35.6 15.40 31.60 40.40 26.03 AnyAttack-Bi w/ Aux 14.6 32.2 44.2 19.00 36.60 45.00 31.93 ViT-L/14 AttackVLM-ii 0.4 1.0 2.6 0.60 1.20 1.80 1.27 AttackVLM-it 0.6 0.8 2.8 0.60 1.20 2.40 1.40 SASD-WS-Cos 1.8 9.2 13.0 3.40 11.40 17.40 9.37 SASD-WS-MSE 3.0 7.8 15.8 4.00 10.60 17.60 9.80 SU-Cos 2.0 8.4 14.2 5.20 13.80 19.60 10.53 SU-MSE 2.0 7.6 12.8 4.00 12.80 15.60 9.13 AnyAttack-Cos 4.6 11.8 18.2 10.60 20.20 25.60 15.17 AnyAttack-Bi 5.4 15.8 21.8 10.20 22.40 29.40 17.50 AnyAttack-Cos w/ Aux 7.8 22.8 33.0 13.80 29.20 38.40 24.16 AnyAttack-Bi w/ Aux 12.4 30.6 41.0 15.20 34.60 44.80 29.77 ViT-L/14 $\times$ 336 AttackVLM-ii 0.4 0.8 2.4 0.40 1.60 2.00 1.27 AttackVLM-it 0.6 1.0 2.4 0.00 1.60 2.20 1.30 SASD-WS-Cos 3.0 8.4 13.8 4.40 14.00 19.20 10.47 SASD-WS-MSE 3.0 9.0 14.4 3.60 11.60 19.00 10.10 SU-Cos 2.6 7.4 10.6 5.20 11.60 15.80 8.87 SU-MSE 3.0 8.0 12.2 5.60 11.80 17.40 9.67 AnyAttack-Cos 4.6 10.0 14.6 8.00 17.40 22.60 12.87 AnyAttack-Bi 4.2 10.8 17.4 8.20 19.00 26.20 14.30 AnyAttack-Cos w/ Aux 8.0 20.2 30.4 13.80 30.40 38.40 21.87 AnyAttack-Bi w/ Aux 9.8 26.0 34.2 16.20 34.60 45.00 27.63
We also report the retrieval performance of our method and the baseline methods on the Flickr30k dataset, with results shown in Table 6. The conclusions are consistent with those obtained from the MSCOCO dataset.
Additional Results on Image Captioning
We present additional examples of the Image Captioning task in Figures 6,  7, and  8. The visualized results further demonstrate the effectiveness of our approach. Interestingly, MiniGPT-4 models (including MiniGPT-v2) tend to generate longer responses. While this behavior slightly affects their quantitative performance on image captioning metrics (as shown in Table 4), the qualitative results suggest that their output remains quite effective.
Figure 6: Visual examples of the image captioning task (Example 1).
Figure 7: Visual examples of the image captioning task (Example 2).
Figure 8: Visual examples of the image captioning task (Example 3).
Additional Results on Commercial VLMs
Figure 9: Responses from three commercial VLMs to our targeted adversarial image (Example 1).
Figure 10: Responses from three commercial VLMs to our targeted adversarial image (Example 2).
Figure 11: Responses from three commercial VLMs to our targeted adversarial image (Example 3).
To further demonstrate the effectiveness of our method, we present additional examples in Figures 9,  10, and  11, where we transfer our attack to commercial VLMs, including Claude’s Sonnet, Microsoft’s Copilot, and Google’s Gemini. The adversarial images generated by our pre-trained decoder were uploaded to these models using their publicly available web interfaces. No prior context or constraints were provided, and the only prompt used was “Describe this image”.

### v2

@@ 4 Experiments
Model Attack Method MSCOCO TR@1 TR@5 TR@10 IR@1 IR@5 IR@10 R@Mean ViT-B/16 AttackVLM-ii 0.4 1.0 1.4 0.24 1.08 2.16 1.05 AttackVLM-it 0.2 1.4 1.8 0.16 1.16 2.12 1.14 SASD-WS-Cos 6.0 17.0 24.8 9.08 24.39 34.55 19.30 SASD-WS-MSE 4.8 18.4 25.6 8.20 25.87 35.15 19.67 SU-Cos 6.8 20.4 27.8 11.11 25.70 33.34 20.86 SU-MSE 6.8 20.6 27.0 10.83 25.10 32.62 20.49 AnyAttack-Cos 8.6 21.2 29.6 10.80 27.59 37.50 22.55 AnyAttack-Bi 12.2 26.2 33.8 12.63 31.71 40.86 26.23 AnyAttack-Cos w/ Aux 8.4 24.8 33.0 11.59 32.10 44.98 25.81 AnyAttack-Bi w/ Aux 14.8 36.8 48.0 17.59 42.02 56.05 35.88 ViT-L/14 AttackVLM-ii 0.2 1.0 1.6 0.24 0.60 1.32 0.83 AttackVLM-it 0.4 0.8 1.4 0.12 0.76 1.48 0.83 SASD-WS-Cos 3.8 11.6 18.8 7.20 18.43 26.35 14.36 SASD-WS-MSE 5.4 14.6 20.6 6.00 18.23 26.47 15.22 SU-Cos 3.0 10.4 13.2 6.19 14.99 20.07 11.31 SU-MSE 3.4 11.2 17.4 6.63 15.27 19.79 12.28 AnyAttack-Cos 3.8 14.0 22.8 7.36 20.71 27.55 16.04 AnyAttack-Bi 4.8 16.0 23.6 8.20 22.31 29.11 17.34 AnyAttack-Cos w/ Aux 9.4 24.6 37.0 11.51 32.62 48.18 27.22 AnyAttack-Bi w/ Aux 12.0 34.0 47.4 15.67 39.34 53.54 33.66 ViT-L/14 $\times$ 336 AttackVLM-ii 0.2 0.6 1.6 0.16 1.12 2.04 0.95 AttackVLM-it 0.2 0.6 1.8 0.32 0.96 1.76 0.94 SASD-WS-Cos 2.8 10.8 16.4 6.52 18.31 26.19 13.50 SASD-WS-MSE 4.4 13.6 19.2 6.72 18.23 25.71 14.64 SU-Cos 2.4 8.0 11.2 4.88 13.79 18.39 9.78 SU-MSE 3.6 8.2 13.2 6.40 14.51 19.19 10.85 AnyAttack-Cos 4.6 11.0 16.6 5.96 17.67 24.23 13.34 AnyAttack-Bi 3.6 14.4 19.0 7.64 19.79 26.83 15.21 AnyAttack-Cos w/ Aux 9.0 23.2 37.2 11.68 34.03 47.62 27.12 AnyAttack-Bi w/ Aux 12.0 33.2 46.8 14.79 39.22 53.06 33.18
@@ 4.1 Experimental Setup
We first employed state-of-the-art targeted adversarial attack on VLMs, AttackVLM [35]. This method includes two variations: AttackVLM-ii and AttackVLM-it, which are based on different attack objectives. Both methods utilize the CLIP ViT-B/32 image encoder as the surrogate model, consistent with our approach. Additionally, we incorporated two targeted adversarial attacks designed for visual classification models: SU [26] and SASD-WS [27]. Since the original cross-entropy loss used in these methods is not suitable for vision-language tasks, we modified them to employ cosine loss and mean squared error (MSE) loss to match targeted images. These modified methods are denoted as SU-Cos/SASD-WS-Cos and SU-MSE/SASD-WS-MSE, respectively. For the SU attack, the surrogate model is CLIP ViT-B/32. For the SASD-WS attack, we utilized the officially released weights, as its surrogate model with the auxiliary model. We denote our proposed methods as AnyAttack-Cos, AnyAttack-Bi, AnyAttack-Cos w/ Aux, and AnyAttack-Bi w/ Aux. These represent AnyAttack fine-tuned with $\mathcal{L}_{\text{Cos}}$, fine-tuned with $\mathcal{L}_{\text{Bi}}$, fine-tuned with $\mathcal{L}_{\text{Cos}}$ using auxiliary models, and fine-tuned with $\mathcal{L}_{\text{Bi}}$ using auxiliary models, respectively.
Datasets, Models, and Tasks.
@@
For the downstream datasets, we utilize the MSCOCO, Flickr30K, and SNLI-VE datasets. We employ a variety of target models, including CLIP, BLIP, BLIP2, InstructBLIP, and MiniGPT-4. The downstream tasks we focus on are image-text retrieval, multimodal classification, and image captioning. For each task, we selected the top 1,000 images. Additionally, following the methodology outlined in [35], we used the top 1,000 images from the ImageNet-1K validation set as clean (random) images to generate adversarial examples.
Figure 3: Example responses from commercial VLMs to targeted attacks generated by our method.
Metric.
@@
We use the attack success rate (ASR) as the primary evaluation metric to assess the performance of targeted adversarial attacks. The calculation of ASR varies slightly depending on the specific task. For instance, in image-text retrieval tasks, ASR represents the recall rate between adversarial images and their corresponding ground-truth text descriptions. In multimodal classification tasks, ASR refers to the accuracy of correctly classifying pairs of “adversarial image and ground-truth description.”
Implementation Details.
@@
In this work, we study perturbations constrained by the $\ell_{\infty}$ norm, ensuring that $\|\delta\|_{\infty}\leq\epsilon$, where $\epsilon$ represents the maximum allowable perturbation, set to $\frac{16}{255}$. We pre-trained the decoder for 520,000 steps on the LAION-400M dataset [20], using a batch size of 600 per GPU on three NVIDIA A100 80GB GPUs. The optimizer used was AdamW, with an initial learning rate of $1\times 10^{-4}$, which was adjusted using cosine annealing. For the downstream datasets, we fine-tuned the decoder for 20 epochs using the same optimizer, initial learning rate, and cosine annealing schedule. We deployed two auxiliary models: the ViT-B/16 trained from scratch on ImageNet-1K and the ViT-L/14 EVA model [6, 7], both of which are trained on ImageNet-1K. The factor $K$ was set to 5. In the pre-training stage, the initial temperature $\tau_{0}$ was set to 1, the final temperature $\tau_{\text{final}}$ was set to 0.07, and the total steps $T$ were set to 10,000. More details can be found in the Appendix.
@@ 4.2 Evaluation on Image-Text Retrieval
In this subsection, we compare the performance of our method against baseline approaches on the image-text retrieval task. Tab. 1 presents the results on the MSCOCO dataset, while results on the Flickr30K dataset are detailed in Appendix. The following key observations can be made:
@@
  • Advantages of Bidirectional Loss: The bidirectional contrastive loss $\mathcal{L}_{\text{Bi}}$ shows clear advantages for retrieval tasks, with AnyAttack-Bi consistently outperforming AnyAttack-Cos.
@@ 4.3 Evaluation on Multimodal Classification
Here, we compare the performance of our attack with the baselines on the multimodal classification task. Tab. 2 presents the results on the SNLI-VE dataset. Our method, AnyAttack-Cos w/ Auxiliary, achieves the highest performance, surpassing the strongest baseline, SASD-WS-MSE, by 20.0%. This underscores the effectiveness of our attack in multimodal classification tasks.
@@ 4.4 Evaluation on Image Captioning
Here, we evaluate the performance of our attack on the image captioning task using the MSCOCO dataset. The VLMs take adversarial images as input and generate text descriptions, which are then assessed against the ground-truth captions using standard metrics. Tab. 3 presents the results across four VLMs: InstructBLIP, BLIP2, BLIP, and MiniGPT-4. Our proposed attack method, AnyAttack-Cos w/ Auxiliary consistently demonstrates superior performance across all evaluation metrics, outperforming the baseline attacks on each VLM.
Attack Method Google Gemini OpenAI GPT AttackVLM-ii 0 2 AttackVLM-it 0 0 SASD-WS-Cos 5 28 SASD-WS-MSE 1 25 SU-Cos 0 11 SU-MSE 0 12 AnyAttack-Cos w/ Aux 31 38
Table 4: Quantitative performance comparison on commercial VLMs. The reported values represent the ASR.
Figure 4: Performance comparison between different configurations of AnyAttack for the image-text retrieval task on MSCOCO. The plot shows the comparative performance of decoder initialized from scratch (Scratch), pre-trained (Pre), and fine-tuned (Cos and Bi), alongside the impact of auxiliary models (w/ Aux) and different fine-tuning objectives (Cos or Bi) on retrieval tasks.
@@ 4.5 Transfer to Commercial VLMs
We evaluate the transferability of adversarial images generated by our method to commercial VLMs: Google Gemini, Claude Sonnet, OpenAI GPT, and Microsoft Copilot. Detailed setups are provided in Appendix.
Quantitative Results.
We selected 100 images from the MSCOCO dataset as target images and conducted a comparison between our method and baseline approaches. Using the prompt “Evaluate the relationship between the given image and text”, along with corresponding options, Tab. 4 reports the percentage of responses labeled as “relevant” by these commercial VLMs. Our method consistently outperforms the baselines, achieving substantial improvements across all evaluated models.
Qualitative Results.
To further demonstrate the effectiveness of AnyAttack, we uploaded the adversarial images to the publicly available web interfaces of these commercial VLMs. Fig. 3 showcases representative examples, with additional instances provided in Appendix. The portions of the VLM responses highlighted in red correspond to the target images, clearly illustrating the impact of our attacks on these VLMs.
@@ 4.6 Further Analysis
Figure 5: Comparison of memory usage and time consumption across different methods.
Ablation Study.
We perform an ablation study on the MSCOCO dataset for image-text retrieval task to evaluate the impact of three key components in our approach: 1) Training approach: Pre-trained, fine-tuned, or trained from scratch. 2) Auxiliary models: With or without auxiliary model integration. 3) Fine-tuning objective: Cosine similarity ($\mathcal{L}_{\text{Cos}}$) vs. bidirectional contrastive loss ($\mathcal{L}_{\text{Bi}}$).
The results, summarized in Fig. 4, reveal the following: 1) Training approach: Fine-tuning a pre-trained model achieves the highest performance, while training from scratch yields significantly worse results, indicating that pre-training is critical for task adaptation. 2) Auxiliary models: The inclusion of auxiliary models consistently improves performance, highlighting their role in enhancing transferability. 3) Fine-tuning objective: The bidirectional contrastive loss ($\mathcal{L}_{\text{Bi}}$) consistently outperforms the cosine similarity loss ($\mathcal{L}_{\text{Cos}}$), demonstrating its effectiveness in improving the alignment of image and text embeddings.
Efficiency Analysis.
@@ 5 Conclusion
In this paper, we introduced AnyAttack, a novel self-supervised framework for generating targeted adversarial attacks on VLMs. Our approach overcomes the scalability limitations of previous methods by enabling the use of any image to serve as a target for attack target without label supervision. Through extensive experiments, we demonstrated the effectiveness of AnyAttack across multiple VLMs and vision-language tasks, revealing significant vulnerabilities in state-of-the-art models. Our method showed considerable transferability, even to commercial VLMs, highlighting the broad implications of our findings.